# RobustVLA: Robustness-Aware Reinforcement Post-Training for Vision-Language-Action Models

### v1

@@ A.3 Proof of robustness theories
Let $d_{t}=\|s_{t}-s_{t}^{*}\|$. Then the state deviation satisfies:
$d_{t+1}\leq L_{f}\cdot d_{t}+L_{f}(\lambda\epsilon_{s}+\epsilon_{t}+\|\xi_{t}\|),$
where $\epsilon_{t}:=\|\pi_{t}-\pi^{*}\|\leq\epsilon_{\text{offline}}+\sum_{i=1}^{t}\delta_{i}.$
@@
Decompose the full error:
@@
Then, by Lipschitz dynamics:
$d_{t+1}=\|f(s_{t},a_{t})-f(s_{t}^{*},a_{t}^{*})\|\leq L_{f}(\|s_{t}-s_{t}^{*}\|+\|a_{t}-a_{t}^{*}\|),$
$\Rightarrow d_{t+1}\leq L_{f}d_{t}+L_{f}(\lambda\epsilon_{s}+\epsilon_{t}+\|\xi_{t}\|).$
@@
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(H^{2}L_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\sum_{t=1}^{H}\delta_{t}+\lambda\epsilon_{s}+\sigma\sqrt{d}\right).$
@@
Define:
$\beta_{t}:=L_{f}(\lambda\epsilon_{s}+\epsilon_{\text{offline}}+\sum_{i=1}^{t}\delta_{i}+\|\xi_{t}\|).$
Then, according to Lemma 4, we have:
$d_{t+1}\leq L_{f}d_{t}+\beta_{t}\quad\Rightarrow\quad d_{t}\leq\sum_{i=0}^{t-1}L_{f}^{t-1-i}\beta_{i}.$
Since $\sum_{i=0}^{t-1}\sum_{j=1}^{i}\delta_{j}=\mathcal{O}(t\sum\delta_{j})$, we have:
$d_{t}\leq\mathcal{O}(tL_{f}^{t})\cdot(\lambda\epsilon_{s}+\epsilon_{\text{offline}}+\sum\delta_{j}+\|\xi_{i}\|).$
At each step:
$|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\leq L_{r}(d_{t}+\lambda\epsilon_{s}+\epsilon_{t}+\|\xi_{t}\|).$
Summing over $H$ steps, and bounding all terms by worst-case geometric growth:
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(H^{2}L_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\sum_{t=1}^{H}\delta_{t}+\lambda\epsilon_{s}+\sigma\sqrt{d}\right).$
@@
Under the setting of Theorem 1, suppose the imitation policy is optimal ($\epsilon_{\text{offline}}=0$), and the dynamics are contractive ($L_{f}<1$). Then the return gap satisfies:
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(H)\cdot L_{r}L_{f}\cdot\lambda\epsilon_{s}.$
@@
From Theorem 1, in the absence of model drift and action noise, we have the recursive state deviation bound:
$d_{t+1}\leq L_{f}d_{t}+L_{f}\lambda\epsilon_{s},$
where $d_{t}:=\|s_{t}-s_{t}^{*}\|$ and $\|\pi(\tilde{s}_{t})-\pi(s_{t})\|\leq\lambda\epsilon_{s}$ due to input perturbation and Jacobian regularization.
Let $B:=L_{f}\lambda\epsilon_{s}$, and unroll the recurrence:
$d_{t}\leq B\sum_{i=0}^{t-1}L_{f}^{i}=B\cdot\frac{1-L_{f}^{t}}{1-L_{f}}\leq\frac{B}{1-L_{f}}.$
Then the return gap is bounded as:
$J(\pi^{*})-J(\pi)=\sum_{t=1}^{H}\left(r(s_{t}^{*},a_{t}^{*})-r(s_{t},a_{t})\right)\leq L_{r}\sum_{t=1}^{H}d_{t}\leq H\cdot L_{r}\cdot\frac{L_{f}\lambda\epsilon_{s}}{1-L_{f}}.$
Therefore, the return gap grows linearly in $H$ and observation noise $\epsilon_{s}$:
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(H)\cdot L_{r}L_{f}\cdot\lambda\epsilon_{s}.$
@@
Under the setting of Theorem 3 but with discounted return $J(\pi)=\mathbb{E}\Big[\sum_{t\geq 1}\gamma^{t-1}r(s_{t},a_{t})\Big]$ for some discount factor $\gamma\in(0,1)$. The dynamics Lipschitz constant and discount satisfy $\gamma L_{f}<1$.The per-step driving term $\zeta_{t}:=\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}+\sum_{i=1}^{t}\delta_{i}+\|\xi_{t}\|$ is uniformly bounded: $\sup_{t}\zeta_{t}\leq C<\infty$. The model deviation sequence $\{\epsilon_{t}\}$ satisfies $\sup_{t}\epsilon_{t}\leq E<\infty$. The per-step update bound $\|\pi_{t}-\pi_{t-1}\|_{\infty}\leq\delta$ with $\delta=o(1/H)$ as $H\to\infty$ (so that, with a good initialization, $\sup_{t}\epsilon_{t}=O(1)$). Then the discounted return gap is bounded independently of horizon $H$:
$J(\pi^{*})-J(\pi)=\mathcal{O}(1).$
@@
Start from the Lipschitz reward difference at time $t$:
$\Delta r_{t}:=\big|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})\big|\leq L_{r}\big(d_{t}+\lambda\epsilon_{s}+\epsilon_{t}+\|\xi_{t}\|\big).$
By Lemma 4, we may unroll the recursion to obtain (for any $t\geq 1$)
$d_{t}\leq L_{f}^{t}d_{0}+L_{f}\sum_{j=0}^{t-1}L_{f}^{\,t-1-j}\zeta_{j}\leq L_{f}^{t}d_{0}+L_{f}C\sum_{j=0}^{t-1}L_{f}^{j}=L_{f}^{t}d_{0}+\frac{L_{f}}{1-L_{f}}\,C.$
Define
$D:=\sup_{t}\big(\lambda\epsilon_{s}+\epsilon_{t}+\mathbb{E}\|\xi_{t}\|\big)<\infty.$
Take the discounted sum of per-step reward differences and apply the bound on $d_{t}$:
$J(\pi^{*})-J(\pi)$ $\leq L_{r}\sum_{t\geq 1}\gamma^{t-1}\big(d_{t}+\lambda\epsilon_{s}+\epsilon_{t}+\mathbb{E}\|\xi_{t}\|\big)$
$\leq L_{r}\sum_{t\geq 1}\gamma^{t-1}\Big(L_{f}^{t}d_{0}+\frac{L_{f}}{1-L_{f}}C\Big)\;+\;L_{r}\sum_{t\geq 1}\gamma^{t-1}D.$
Evaluate each geometric series separately. For the term involving $L_{f}^{t}d_{0}$,
$\sum_{t\geq 1}\gamma^{t-1}L_{f}^{t}d_{0}=d_{0}L_{f}\sum_{t\geq 1}(\gamma L_{f})^{t-1}=d_{0}\frac{L_{f}}{1-\gamma L_{f}},$
which is finite by the assumption $\gamma L_{f}<1$. For the constant term,
$\sum_{t\geq 1}\gamma^{t-1}\frac{L_{f}}{1-L_{f}}C=\frac{L_{f}}{1-L_{f}}C\sum_{t\geq 1}\gamma^{t-1}=\frac{L_{f}}{1-L_{f}}C\cdot\frac{1}{1-\gamma}.$
Finally,
$\sum_{t\geq 1}\gamma^{t-1}D=\frac{D}{1-\gamma}.$
Combining these bounds yields
$J(\pi^{*})-J(\pi)\leq L_{r}\left(d_{0}\frac{L_{f}}{1-\gamma L_{f}}+\frac{L_{f}}{1-L_{f}}\,\frac{C}{1-\gamma}+\frac{D}{1-\gamma}\right),$
which is a finite constant independent of the planning horizon $H$. Thus, we obtain:
$J(\pi^{*})-J(\pi)=\mathcal{O}(1).$
@@ A.4 Observation perturbation setting
Following previous work, GEVRM (Zhang et al., 2025a), we provide details on implementing observation perturbations in the LIBERO simulation platform: 1) Image Shift: The image state is randomly shifted to the upper left, with a maximum translation rate of $0.3$ relative to the image size. 2) Image Rotation: The image state is randomly rotated counterclockwise with a maximum rotation angle of $30$ degrees. 3) Color Jitter: The image state saturation, brightness, contrast, and sharpness are randomly jittered with a maximum random factor of $3$. 4) Image Occlusions: The image state is randomly occluded with a random number of occlusion blocks ranging from $1$ to $3$ and a maximum length of $20$. 5) Image Erasing: The image state is perturbed by random noise blocks with a maximum size of $0.1$ of the original image. We add observation perturbations to both the first-view and third-view observation images, which makes the constructed robust benchmark platform more challenging.
@@ A.5 Implementation Details
Reward Densification. Given that only sparse 0-1 rewards representing task success or failure are directly obtained from the environment, we follow the previous work, ReinboT (Zhang et al., 2025c), to densify the rewards. Specifically, in addition to the task completion reward obtained from the environment, we consider ten additional rewards, as shown in Tab. 7. The ORB-related reward from ReinboT is not considered, as its computation is somewhat time-consuming during online interaction.
Experimental implementation. We conduct our experiments on OpenVLA-OFT (Kim et al., 2025) using the official checkpoints as the base model, with RIPT-VLA’s LoRA adaptor checkpoints (Tan et al., 2025b) applied for adaptation. The implementation is based on the RIPT-VLA codebase (Tan et al., 2025b). Moreover, $\mathcal{R}_{\text{Jac}}$ requires computing the two-norm of $\nabla_{s}\log\pi_{\theta}(a|s)$, which is expensive and memory-intensive to compute directly at the image pixels. Thus, we compute the Jacobian on the low-dimensional embeddings obtained from the Llama-2 encoding used by OpenVLA-OFT. Training is performed on a single GPU with LoRA of rank 32. The post-training procedure takes approximately 24 hours for fine-tuning. The training parameters are shown in Tab. 6, and other parameters follow the previous work, RIPT-VLA (Tan et al., 2025b). All experiments are conducted on the following hardware: CPU: Intel(R) Xeon(R) Platinum 8358 @ 2.60GHz; GPU: NVIDIA A100-SXM4-80GB.
Baseline algorithm reproduction. To replicate the baseline algorithms, $\pi_{0}$, ReinboT, RWR, and ARFM are all implemented in the repositories: huggingface/lerobot (Cadene et al., 2024). To replicate GEVRM, we utilize the same prototype contrastive learning method for observation representation in the RIPT-VLA code. For other baseline models, we utilize the official implementation.

### v2

@@ A.3 Proof of robustness theories
Let $d_{t}=\|s_{t}-s_{t}^{*}\|$. Under $L_{f}$-Lipschitz dynamics, and with both observation perturbations $\|\tilde{s}_{t}-s_{t}\|\leq\epsilon_{s}$ and action noise $\xi_{t}$, the state deviation satisfies
$d_{t+1}\leq L_{f}\cdot d_{t}+L_{f}\big(\lambda\epsilon_{s}+\epsilon_{t}+\|\xi_{t}\|\big),$
where, for all $t$, $\epsilon_{t}:=\|\pi_{t}-\pi^{*}\|\leq\epsilon_{\mathrm{offline}}+{\sum_{i=1}^{N}\delta_{i}.}$
@@
Decompose the action error at time $t$:
@@
Applying $L_{f}$-Lipschitz dynamics,
$d_{t+1}=\|f(s_{t},a_{t})-f(s_{t}^{*},a_{t}^{*})\|\leq L_{f}\big(\|s_{t}-s_{t}^{*}\|+\|a_{t}-a_{t}^{*}\|\big),$
which yields the stated recursion. ∎
@@
$\mathbb{E}[J(\pi^{*})-J(\pi)]\;\leq\;\mathcal{O}(HL_{r}L_{f}^{H})\,\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\lambda\epsilon_{s}+\sigma\sqrt{d}\Big).$}
@@
From Lemma 4, the state deviation satisfies
$d_{t+1}\;\leq\;L_{f}d_{t}+L_{f}\big(\lambda\epsilon_{s}+\epsilon_{t}+\|\xi_{t}\|\big),\quad\epsilon_{t}\;\leq\;\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}.$
Unrolling the linear recurrence with $d_{0}=0$ yields
$d_{t}\;\leq\;\sum_{j=0}^{t-1}L_{f}^{\,t-1-j}\,\big(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\|\xi_{j}\|\big).$
Taking expectations and using linearity,
$\mathbb{E}[d_{t}]\;\leq\;\sum_{j=0}^{t-1}L_{f}^{\,t-1-j}\Big(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\mathbb{E}\|\xi_{j}\|\Big).$
Applying Lemma 3, $\mathbb{E}\|\xi_{j}\|\leq\sigma\sqrt{d}$, gives
$\mathbb{E}[d_{t}]\;\leq\;\Big(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\sigma\sqrt{d}\Big)\sum_{j=0}^{t-1}L_{f}^{t-1-j}.$
Thus,
$\mathbb{E}[d_{t}]\;\leq\;\mathcal{O}(L_{f}^{t})\Big(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\sigma\sqrt{d}\Big).$
Because rewards are $L_{r}$-Lipschitz,
$|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\;\leq\;L_{r}\big(d_{t}+\lambda\epsilon_{s}+\epsilon_{t}+\|\xi_{t}\|\big).$
Similar to the proof of Theorem 2, we obtain:
$\mathbb{E}[J(\pi^{*})-J(\pi)]\;\leq\;\mathcal{O}(HL_{r}L_{f}^{H})\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\lambda\epsilon_{s}+\sigma\sqrt{d}\Big).$
Proof of Corollaries 1 and 2
@@
Under the assumptions of Theorem 1 but with $\epsilon_{\mathrm{offline}}=0$ (the imitation model is perfect) and contractive dynamics $L_{f}<1$, the expected return gap satisfies
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\;\leq\;H\cdot L_{r}\cdot\frac{1}{1-L_{f}}\cdot\lambda\epsilon_{s}.$}
@@
With $\epsilon_{\mathrm{offline}}=0$, Lemma 1 gives (for $d_{0}=0$)
$d_{t}\leq L_{f}(\lambda\epsilon_{s})\sum_{i=0}^{t-1}L_{f}^{\,i}=\lambda\epsilon_{s}\cdot L_{f}\cdot\frac{1-L_{f}^{\,t}}{1-L_{f}}.$
Since $L_{f}\in(0,1)$, the factor $L_{f}\frac{1-L_{f}^{\,t}}{1-L_{f}}\leq\frac{L_{f}}{1-L_{f}}$ uniformly in $t$. Thus
$d_{t}\leq\lambda\epsilon_{s}\cdot\frac{L_{f}}{1-L_{f}}.$
The per-step reward difference satisfies
$|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\leq L_{r}\big(d_{t}+\lambda\epsilon_{s}\big)\leq L_{r}\lambda\epsilon_{s}\Big(\frac{L_{f}}{1-L_{f}}+1\Big)=L_{r}\lambda\epsilon_{s}\cdot\frac{1}{1-L_{f}}.$
Taking expectations and summing over $t=1,\dots,H$ yields
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\leq H\cdot L_{r}\lambda\epsilon_{s}\cdot\frac{1}{1-L_{f}},$
which is the stated bound. ∎
@@
Under the setting of Theorem 3 but with discounted return $J(\pi)=\mathbb{E}\Big[\sum_{t\geq 1}\gamma^{t-1}r(s_{t},a_{t})\Big]$, for some discount factor $\gamma\in(0,1)$ with contractive dynamics $L_{f}\in(0,1)$. Assume the per-step driving term $\zeta_{t}:=\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\|\xi_{t}\|$ is uniformly bounded in expectation, i.e. there exists $C<\infty$ such that $\sup_{t}\mathbb{E}[\zeta_{t}]\leq C$, and that $\sup_{t}\mathbb{E}[\epsilon_{t}]\leq E<\infty$. Then the discounted expected return gap is bounded independently of the rollout horizon $H$:
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]=\mathcal{O}(1).$
@@
From Lemma 4 we have the one-step recursion
$d_{t+1}\leq L_{f}d_{t}+L_{f}\,\zeta_{t},$
where $d_{t}=\|s_{t}-s_{t}^{*}\|$. Unrolling with $d_{0}=0$ gives
$d_{t}\leq\sum_{j=0}^{t-1}L_{f}^{\,t-1-j}\,L_{f}\,\zeta_{j}=\sum_{j=0}^{t-1}L_{f}^{\,t-j}\,\zeta_{j}.$
Taking expectations and using $\mathbb{E}[\zeta_{j}]\leq C$ for all $j$ yields
$\mathbb{E}[d_{t}]\leq C\sum_{j=0}^{t-1}L_{f}^{\,t-j}=C\sum_{k=1}^{t}L_{f}^{\,k}=C\cdot\frac{L_{f}(1-L_{f}^{\,t})}{1-L_{f}}\leq C\cdot\frac{L_{f}}{1-L_{f}},$
so $\mathbb{E}[d_{t}]$ is uniformly bounded in $t$.
The per-step expected reward difference satisfies (by $L_{r}$-Lipschitzness of $r$ and the definition of $\zeta_{t}$)
$\mathbb{E}\big[|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\big]\leq L_{r}\big(\mathbb{E}[d_{t}]+\mathbb{E}[\zeta_{t}]\big)\leq L_{r}\left(C\cdot\frac{L_{f}}{1-L_{f}}+C\right)=L_{r}C\left(\frac{L_{f}}{1-L_{f}}+1\right).$
Therefore
$\mathbb{E}[J(\pi^{*})-J(\pi)]$ $\leq L_{r}\sum_{t\geq 1}\gamma^{t-1}\Big(\mathbb{E}[d_{t}]+\mathbb{E}[\zeta_{t}]\Big)$
$\leq L_{r}C\sum_{t\geq 1}\gamma^{t-1}\left(\frac{L_{f}(1-L_{f}^{\,t})}{1-L_{f}}+1\right).$
Evaluate the geometric series (using $\gamma L_{f}<1$):
$\sum_{t\geq 1}\gamma^{t-1}\frac{L_{f}(1-L_{f}^{\,t})}{1-L_{f}}=\frac{L_{f}}{1-L_{f}}\sum_{t\geq 1}\big(\gamma^{t-1}-(\gamma L_{f})^{t-1}\big)=\frac{L_{f}}{1-L_{f}}\Big(\frac{1}{1-\gamma}-\frac{1}{1-\gamma L_{f}}\Big),$
and
$\sum_{t\geq 1}\gamma^{t-1}=\frac{1}{1-\gamma}.$
Therefore
$\mathbb{E}[J(\pi^{*})-J(\pi)]\leq L_{r}C\left(\frac{L_{f}}{1-L_{f}}\Big(\frac{1}{1-\gamma}-\frac{1}{1-\gamma L_{f}}\Big)+\frac{1}{1-\gamma}\right),$
which is a finite constant independent of the rollout horizon $H$. Thus, we obtain:
$\mathbb{E}[J(\pi^{*})-J(\pi)]=\mathcal{O}(1).$
@@ A.4 Observation perturbation setting
Following previous work, GEVRM (zhang2025gevrm), we provide details on implementing observation perturbations in the LIBERO simulation platform: 1) Image Shift: The image state is randomly shifted to the upper left, with a maximum translation rate of $0.3$ relative to the image size. 2) Image Rotation: The image state is randomly rotated counterclockwise with a maximum rotation angle of $30$ degrees. 3) Color Jitter: The image state saturation, brightness, contrast, and sharpness are randomly jittered with a maximum random factor of $3$. 4) Image Occlusions: The image state is randomly occluded with a random number of occlusion blocks ranging from $1$ to $3$ and a maximum length of $20$. 5) Image Erasing: The image state is perturbed by random noise blocks with a maximum size of $0.1$ of the original image. We add observation perturbations to both the first-view and third-view observation images, which makes the constructed robust benchmark platform more challenging.
@@ A.5 Implementation Details
Reward Densification. Given that only sparse 0-1 rewards representing task success or failure are directly obtained from the environment, we follow the previous work, ReinboT (zhang2025reinbot), to densify the rewards. Specifically, in addition to the task completion reward obtained from the environment, we consider ten additional rewards, as shown in Tab. 7. The ORB-related reward from ReinboT is not considered, as its computation is somewhat time-consuming during online interaction.
Experimental implementation. We conduct our experiments on OpenVLA-OFT (kim2025fine) using the official checkpoints as the base model, with RIPT-VLA’s LoRA adaptor checkpoints (tan2025interactiveposttrainingvisionlanguageactionmodels) applied for adaptation. The implementation is based on the RIPT-VLA codebase (tan2025interactiveposttrainingvisionlanguageactionmodels). Moreover, $\mathcal{R}_{\text{Jac}}$ requires computing the two-norm of $\nabla_{s}\log\pi_{\theta}(a|s)$, which is expensive and memory-intensive to compute directly at the image pixels. Thus, we compute the Jacobian on the low-dimensional embeddings obtained from the Llama-2 encoding used by OpenVLA-OFT. Training is performed on a single GPU with LoRA of rank 32. The post-training procedure takes approximately 24 hours for fine-tuning. The training parameters are shown in Tab. 6, and other parameters follow the previous work, RIPT-VLA (tan2025interactiveposttrainingvisionlanguageactionmodels). All experiments are conducted on the following hardware: CPU: Intel(R) Xeon(R) Platinum 8358 @ 2.60GHz; GPU: NVIDIA A100-SXM4-80GB.
Baseline algorithm reproduction. To replicate the baseline algorithms, $\pi_{0}$, ReinboT, RWR, and ARFM are all implemented in the repositories: huggingface/lerobot (cadene2024lerobot). To replicate GEVRM, we utilize the same prototype contrastive learning method for observation representation in the RIPT-VLA code. For other baseline models, we utilize the official implementation.
@@ A.6 Use of Large Language Models
Figure 7: Action distribution during model inference under four noise conditions.
Figure 8: Comparison of action distribution between baseline RIPT-VLA and the proposed RobustVLA.